# Xingwang Li, , Dusit Niyato, and Chau Yuen

### v1

@@ Abstract
Flexible intelligent metasurfaces (FIMs) offer a new solution for wireless communications by introducing morphological degrees of freedom, dynamically morphing their three-dimensional shape to ensure multipath signals interfere constructively. However, realizing the desired performance gains in FIM systems is critically dependent on acquiring accurate channel state information across a continuous and high-dimensional deformation space. Therefore, this paper investigates this fundamental channel estimation problem for FIM assisted millimeter-wave communication systems. First, we develop model-based frameworks that structure the problem as either function approximation using interpolation and kernel methods or as a sparse signal recovery problem that leverages the inherent angular sparsity of millimeter-wave channels. To further advance the estimation capability beyond explicit assumptions in model-based channel estimation frameworks, we propose a deep learning-based framework using a Fourier neural operator (FNO). By parameterizing a global convolution operator in the Fourier domain, we design an efficient FNO architecture to learn the continuous operator that maps FIM shapes to channel responses with mesh-independent properties. Furthermore, we exploit a hierarchical FNO (H-FNO) architecture to efficiently capture the multi-scale features across a hierarchy of spatial resolutions. Numerical results demonstrate that the proposed H-FNO significantly outperforms the model-based benchmarks in estimation accuracy and pilot efficiency. In particular, the interpretability analysis show that the proposed H-FNO learns an anisotropic spatial filter adapted to the physical geometry of FIM and is capable of accurately reconstructing the non-linear channel response across the continuous deformation space.
@@ I Introduction
The relentless pursuit of higher data rates, ubiquitous connectivity, and improved energy efficiency for the sixth-generation (6G) wireless networks has spurred the exploration of novel technologies that manipulate the wireless propagation environment itself [1]. Metasurfaces and other reconfigurable array technologies have garnered significant attention by offering the ability to passively reflect and steer electromagnetic waves, thereby creating favorable channel conditions with reduced hardware complexity [2]. However, these technologies are predominantly based on rigid planar structures, e.g., reconfigurable intelligent surfaces (RISs), which limits their ability to adapt to complex propagation environments [3, 4, 5], especially when deep fading occurs due to destructive multipath interference. Flexible intelligent metasurfaces (FIMs) introduce a revolutionary paradigm by endowing the transceiver with physical adaptability [6]. An FIM is a conformal array whose radiating elements can be precisely displaced, typically in a direction perpendicular to the surface, allowing the entire structure to morph into a non-planar three-dimensional (3D) shape [7]. This additional morphological degree of freedom allows the FIM to actively reconfigure its geometry to ensure that multipath signal components add constructively, thus providing a powerful new mechanism to enhance communication performance beyond what is achievable with rigid arrays [8].
@@ I-A1 Hardware Design of FIM
A critical enabler for the FIMs discussed in this paper is the ability to dynamically and precisely control the surface shape. A key advancement in this area was the development of soft and shape-programmable surfaces actuated by fast electromagnetic control of liquid metal networks [16]. The authors in [17] demonstrated a dynamically reprogrammable surface capable of self-evolving shape morphing [17]. More recently, these concepts have been extended into the microwave regime with FIMs that feature shape-guided adaptive programming [7]. These collective advancements in materials, fabrication, and particularly in dynamic actuation, provide the physical foundation for the FIM hardware.
@@ I-A2 Communication Design of FIM
Complementing the advancements in hardware, a significant body of research has focused on the communication theory and signal processing algorithms required to unlock the potential of FIMs [8]. In fact, the underlying principle of leveraging physical deformation for communication enhancement has been preliminarily explored through related concepts, such as movable antennas [18], fluid antennas [19, 20], and pinching-antenna systems [21, 22], which have demonstrated that changing antenna positions can yield significant selection diversity gains, particularly in high-frequency bands. Building upon this principle of flexible arrays, FIMs advance this principle by integrating the 3D flexibility with the large aperture and element density characteristic of metasurfaces.
At present, research into FIM-aided communication systems is still in its nascent stages. Initial representative works have begun to establish the potential performance gains across different scenarios. In [9], the joint optimization of transmit beamforming and the FIM surface shape has been investigated to proactively manage inter-user interference and enhance multiuser performance for FIM-aided multiuser systems. Furthermore, the authors in [10] investigated FIM assisted MIMO systems, where the joint optimization of the FIM 3D surface shape and the transmit covariance matrix was proposed to maximize channel capacity. In [23], the general models for flexible antenna arrays were established and evaluated, which consider various transformations e.g., bending, folding, and rotation, to characterize the resulting channel variations. Beyond enhancing data rates, the unique capabilities of FIMs have been explored for other applications, most notably wireless sensing. the authors in [11] investigated the joint design of FIM surface shapes and transmit waveforms to enhance multi-target sensing performance.
@@ I-B Motivations and Contributions
A common thread throughout the aforementioned application-focused works is the general assumption that the accurate channel is available. However, acquiring the accurate channel in the continuous and high-dimensional deformation space of FIMs is a fundamental and non-trivial challenge [8]. This motivates our work to develop novel and efficient channel estimation techniques specifically tailored for FIM-aided communication systems. Specifically, unlike rigid arrays where the antenna geometry is fixed, the channel in an FIM system is a function of the high-dimensional deformation vector that represents the displacement of all radiating elements [9, 10]. The space of possible deformations is continuous and infinite, making it impossible to measure the channel for every possible shape. A naive approach of discretizing this high-dimensional space would lead to an intractable number of pilot transmissions, resulting in prohibitive overhead. Therefore, a core challenge is to develop low-overhead channel estimation schemes that can accurately characterize the channel for the entire continuous deformation space based on a limited number of pilot measurements.
@@ I-C Organizations and Notations
II System Model and Problem Formulation
Figure 1: FIM assisted multi-user mmWave communications.
As shown in Fig. 1, we consider downlink communication system where a base station (BS) equipped with an FIM serves $K$ single-antenna users. The FIM consists of $N=N_{x}\times N_{z}$ radiating elements arranged as a uniform planar array (UPA) on a flexible substrate. With the aid of controller that can be driven by micro-electromechanical systems (MEMS) or electromagnetic Lorentz force [6, 17, 16], the FIM dynamically alter its physical shape with programmer manner is harnessed to optimize the propagation channels for all users.
II-A FIM Geometry and Deformation
To describe an FIM with arbitrary orientation in a global coordinate system, we define a local coordinate system for the FIM. This local system is characterized by three angles: azimuth $\phi_{t}\in[0,\pi)$, elevation $\theta_{t}\in[0,\pi)$, and spin $\rho_{t}\in[0,2\pi)$. These angles define the local orthonormal basis vectors $\{\mathbf{i}_{t},\mathbf{j}_{t},\mathbf{k}_{t}\}$ of FIM, where $\mathbf{k}_{t}$ is the normal vector to the undeformed surface, and $\{\mathbf{i}_{t},\mathbf{j}_{t}\}$ are orthogonal vectors spanning the surface plane. Following the model in [10], the basis vectors are given by
@@ II-A FIM Geometry and Deformation
The position of the $n$-th element on the undeformed surfac $\mathbf{p}_{n}$ is defined relative to the center $\mathbf{p}_{c}$ of the FIM, by its local two-dimensional (2D) coordinates $(x_{n}^{t},y_{n}^{t})$:
@@
The core feature of the FIM is the controllable deformation. Each element $n$ can be displaced by a distance $\zeta_{n}\in[-\bar{\zeta},\bar{\zeta}]$ along the normal vector $\mathbf{k}_{t}$, where $\bar{\zeta}>0$ denotes the maximum range allowed by the unilateral reversible deformation of the FIM. The collection of these displacements forms the deformation vector $\bm{\zeta}=[\zeta_{1},\dots,\zeta_{N}]^{T}\in\mathbb{C}^{N\times 1}$. Final, the post-deformation position of the $n$-th element is given by
@@ II-B Channel Model
The channel between the $N$-element FIM at the BS and each of the $K$ users is distinct and depends on the deformation shape $\bm{\zeta}$ of FIM. We adopt a clustered multipath channel model, where for each user $k\in\{1,\dots,K\}$, the channel is formed by $L_{k}$ scattering clusters, and the $l$-th cluster contributes $G_{k,l}$ individual propagation paths. Following the mmWave channel modeling in 3GPP [24], the channel $\mathbf{h}_{k}(\bm{\zeta})\in\mathbb{C}^{N\times 1}$ between BS and user $k$ is given by
@@
where $\mathbf{o}_{k,l,g}$ is the unit-norm direction of departure (AoD) vector for the $g$-th path in the $l$-th cluster to user $k$. ${\mathbf{a}}(\bm{\zeta},\mathbf{o}_{k,l,g})\in\mathbb{C}^{N\times 1}$ is the steering vector of the FIM, which is a function of the deformation $\bm{\zeta}$ and a given AoD $\mathbf{o}$. The AoDs for paths within a single cluster are distributed around a mean cluster angle with a certain angular spread. $\alpha_{k,l,g}$ is the complex gain of the corresponding path, which is defined as [24]
@@
where $d_{k}$ is the distance between the BS and user $k$. $\beta(d_{k})=P_{0}\left(\frac{d_{k}}{d_{0}}\right)^{-\nu}$ is the large-scale path loss coefficient. Here, $P_{0}$ is the path loss at a reference distance $d_{0}$, and $\nu$ is the path loss exponent. $g_{k,l,g}\sim\mathcal{CN}(0,1)$ is the small-scale fading coefficient, modeled as a complex Gaussian random variable.
In FIM systems, the steering vector ${\mathbf{a}}$ can be decomposed into two parts: a rigid component $\mathbf{a}_{\text{rigid}}$ and a flexible component $\mathbf{f}_{\text{flex}}$, i.e.,
@@
$\begin{split}[\mathbf{a}_{\text{rigid}}(\mathbf{o})]_{n}&=e^{\left(j\kappa\langle\mathbf{p}_{n}-\mathbf{p}_{c},\mathbf{o}\rangle\right)}\\ &=e^{\left(j\kappa\langle x_{n}^{t}\mathbf{i}_{t}+y_{n}^{t}\mathbf{j}_{t},\mathbf{o}\rangle\right)},\end{split}$ (7)
@@
$[\mathbf{f}_{\text{flex}}(\bm{\zeta},\mathbf{o})]_{n}=\exp\left(j\kappa\langle\zeta_{n}\mathbf{k}_{t},\mathbf{o}\rangle\right).$ (8)
@@
[…]
@@
Substituting the coordinates, the steering vector in FIM systems becomes
$\begin{split}&\left[\mathbf{a}(\bm{\zeta},\theta_{k,l,g},\phi_{k,l,g})\right]_{n}=\\ &e^{\left(j\frac{2\pi}{\lambda}\left(x_{n}\sin\theta_{k,l,g}\cos\phi_{k,l,g}+\zeta_{n}\sin\theta_{k,l,g}\sin\phi_{k,l,g}+z_{n}\cos\theta_{k,l,g}\right)\right)}.\end{split}$ (11)
@@ II-C Problem Formulation
$\begin{split}{\mathbf{y}_{q}}&=\sum\limits_{k=1}^{K}{\mathbf{{h}}_{k}(\bm{\zeta}_{q}){{{{s}}}_{k,q}}+{\mathbf{w}_{q}}},\end{split}$
@@
Figure 2: Frame structure for FIM channel estimation.
As illustrated in Fig. 2, suppose the uplink pilot transmission is divided $M$ subframes, in which each subframe is composed of $P$ pilot slots. In the $m$-th subframe, the collected signal $\mathbf{Y}_{m}\in{{\mathbb{C}}^{N\times P}}$ at the BS can be expressed as
@@ III-A Interpolation-Based Channel Estimation
[…]
@@
In a general $\gamma$-dimensional space, this concept is extended through barycentric coordinates. First, a simplex, e.g., a triangle in 2D or a tetrahedron in 3D, of $\gamma+1$ pilot points $\{\bm{\zeta}_{1},\dots,\bm{\zeta}_{\gamma+1}\}$ that encloses the target shape $\bm{\zeta}_{\text{target}}$ is found via Delaunay triangulation [26]. The target shape can then be uniquely expressed as a linear combination of the simplex vertices:
$\begin{split}&\bm{\zeta}_{\text{target}}=\sum_{m=1}^{\gamma+1}\lambda_{m}\bm{\zeta}_{m},\\ &\text{s.t.}\quad\sum_{m=1}^{\gamma+1}\lambda_{m}=1\text{ and }\lambda_{m}\geq 0,\end{split}$ (18)
@@
where the coefficients $\{\lambda_{m}\}$ are the barycentric coordinates and serve as the weights in the general interpolation framework. If $\bm{\zeta}_{m}$ is a vertex of the enclosing simplex, $w_{m}(\bm{\zeta}_{\text{target}})=\lambda_{m}$, otherwise $w_{m}(\bm{\zeta}_{\text{target}})=0$. The interpolated channel is then simply $\hat{\mathbf{h}}_{\text{interp}}(\bm{\zeta}_{\text{target}})=\sum_{m=1}^{\gamma+1}\lambda_{m}\hat{\mathbf{h}}(\bm{\zeta}_{m})$. However, this approach faces significant practical challenges in the high-dimensional space of FIM deformations. For an FIM with $N$ antennas, the deformation space is $N$-dimensional. Constructing a single simplex requires at least $N+1$ pilot shapes. This requirement, $M\geq N+1$, often makes rigorous linear or cubic interpolation computationally infeasible for FIMs where the number of pilot measurements $M$ is typically much smaller than the number of antennas $N$. The infeasibility of performing a rigorous Delaunay triangulation in the high-dimensional FIM deformation space motivates a more practical and local approximation of linear interpolation. Instead of constructing a global triangulation, we propose a method based on finding a local linear representation of the target shape using its KNN from the set of pilot shapes [27]. In the KNN estimator, for the specific target shape $\bm{\zeta}_{\text{target}}$, the corresponding channel vector is regarded as a linear combination of the channel vectors of its $\bar{k}$ closest pilot shapes. Specifically, the hyperparameter $\bar{k}$ is chosen to be a small integer satisfying $\bar{k}\ll N+1$. We seek a weight vector $\mathbf{w}=[w_{1},\dots,w_{\bar{k}}]^{T}\in\mathbb{R}^{\bar{k}\times 1}$ that best reconstructs the target shape from its neighbors. This is formulated as a least-squares problem:
$\min_{\mathbf{w}}\quad\left\|\bm{\zeta}_{\text{target}}-\sum_{i=1}^{\bar{k}}w_{i}\bm{\zeta}_{m_{i}}\right\|_{2}^{2}.$ (19)
Let $\bar{\bm{\zeta}}_{\bar{k}}=[\bm{\zeta}_{m_{1}},\dots,\bm{\zeta}_{m_{\bar{k}}}]\in\mathbb{R}^{N\times\bar{k}}$ be the matrix of the $\bar{k}$ nearest neighbor shapes. The above problem can be written as $\min_{\mathbf{w}}\|\bm{\zeta}_{\text{target}}-\bar{\bm{\zeta}}_{\bar{k}}\mathbf{w}\|_{2}^{2}$. The optimal linear weight vector $\mathbf{w}^{*}=[{w}^{*}_{1},\ldots,{w}^{*}_{\bar{k}}]$ for representing the target shape within its local neighborhood that solves this problem is found via
$\mathbf{w}^{*}=(\bar{\bm{\zeta}}_{\bar{k}}^{T}\bar{\bm{\zeta}}_{\bar{k}})^{-1}\bar{\bm{\zeta}}_{\bar{k}}^{T}\bm{\zeta}_{\text{target}}.$ (20)
Finally, the obtained weights are applied to the corresponding noisy channel measurements of the neighbors $\{\hat{\mathbf{h}}({\bm{\zeta}}_{m_{1}}),\dots,\hat{\mathbf{h}}({\bm{\zeta}}_{m_{\bar{k}}})\}$, to synthesize the channel estimate as
$\hat{\mathbf{h}}_{\text{KNN}}({\bm{\zeta}}_{\text{target}})=\sum_{i=1}^{\bar{k}}w_{i}^{*}\hat{\mathbf{h}}({\bm{\zeta}}_{m_{i}}).$ (21)
The KNN-based local linear approach elegantly bypasses the curse of dimensionality associated with global triangulation. By operating in a small, local neighborhood, it remains computationally efficient and numerically stable, while preserving the core principle of linear combination that underlies interpolation methods.

### v2

@@ Abstract
Flexible intelligent metasurfaces (FIMs) offer a new solution for wireless communications by introducing morphological degrees of freedom, dynamically morphing their three-dimensional shape to ensure multipath signals interfere constructively. However, realizing the desired performance gains in FIM systems is critically dependent on acquiring accurate channel state information across a continuous and high-dimensional deformation space. Therefore, this paper investigates this fundamental channel estimation problem for FIM assisted millimeter-wave communication systems. First, we develop model-based frameworks that structure the problem as either function approximation using interpolation and kernel methods or as a sparse signal recovery problem that leverages the inherent sparsity of millimeter-wave channels. To further improve the estimation capability beyond explicit assumptions in model-based channel estimation frameworks, we propose a deep learning-based framework using a Fourier neural operator (FNO). By parameterizing a global convolution operator in the Fourier domain, we design an efficient FNO architecture to learn the continuous operator that maps FIM shapes to channel responses with mesh-independent properties. Furthermore, we exploit a hierarchical FNO (H-FNO) architecture to efficiently capture the multi-scale features across a hierarchy of spatial resolutions. Numerical results demonstrate that the proposed H-FNO significantly outperforms the model-based benchmarks in estimation accuracy and pilot efficiency, while also showing superior robustness to hardware imperfections. In particular, interpretability analysis demonstrates that the proposed H-FNO learns an anisotropic spatial filter adapted to the physical geometry of FIM and decomposes the channel into multi-scale features.
@@ I Introduction
The relentless pursuit of higher data rates, ubiquitous connectivity, and improved energy efficiency for the sixth-generation (6G) wireless networks has spurred the exploration of novel technologies that manipulate the wireless propagation environment itself [1]. Reconfigurable metasurface and other flexible array technologies have garnered significant attention by offering the ability to passively reflect and steer electromagnetic waves, thereby creating favorable channel conditions with reduced hardware complexity [2, 3]. However, these technologies are predominantly based on rigid planar structures, which limits their ability to adapt to complex propagation environments [4, 5, 6], especially when deep fading occurs due to destructive multipath interference. Flexible intelligent metasurfaces (FIMs) introduce a revolutionary paradigm by endowing the transceiver with physical adaptability [7]. An FIM is a conformal array whose radiating elements can be precisely displaced, typically in a direction perpendicular to the surface, allowing the entire structure to morph into a non-planar three-dimensional (3D) shape [8]. This additional morphological degree of freedom allows the FIM to actively reconfigure its geometry to ensure that multipath signal components add constructively, thus providing a powerful new mechanism to enhance communication performance beyond what is achievable with rigid arrays [9].
@@ I-A1 Hardware Design of FIM
A critical enabler for the FIM discussed in this paper is the ability to dynamically and precisely control the surface shape. The dynamic morphing capability of FIM is enabled by integrating a grid of micro-actuators into the flexible substrate that houses the radiating elements, in which a central controller sends signals to this actuator grid, allowing for the rapid adjustment of the metasurface into a desired 3D shape. In [17], a liquid metal architecture that relies on the electromagnetic Lorentz force was proposed to implement precise and low-power element displacement of FIM. The authors in [18] demonstrated a dynamically reprogrammable surface capable of self-evolving shape morphing [18]. More recently, these concepts have been extended into the microwave regime with FIMs that feature shape-guided adaptive programming [8]. These collective advancements in materials, fabrication, and particularly in dynamic actuation, provide the physical foundation for the FIM hardware.
@@ I-A2 Communication Design of FIM
Complementing the advancements in hardware, a significant body of research has focused on the signal processing required to unlock the potential of reconfigurable antenna arrays. These technologies follow two main paradigms: electronic tuning of electromagnetic properties, exemplified by the emerging reconfigurable intelligent surfaces (RISs), and physical reconfiguration of antenna locations, e.g., fluid, movable, and pinching-antenna systems [19, 20, 21]. Despite their different mechanisms, both share the critical challenge of channel estimation due to their massive configurable degrees of freedom. Consequently, research has yielded both model-based and data-driven solutions to address this challenge. Model-based approaches exploit physical priors, leveraging specific channel estimation approaches, e.g., sparsity-based compressed sensing and tensor decomposition [22, 23, 24, 25, 26]. For instance, the authors of [25] proposed a tensor decomposition-based effective channel estimation framework for movable antenna systems. In [26], the authors developed a highly sophisticated polar-domain dictionary to accurately reconstruct the near-field channel with wideband beam split effect. Conversely, data-driven methods leveraged neural networks to learn the complex and non-linear channel mappings by employing various deep learning architectures [27, 28, 29].
Building upon this principle of reconfigurable antennas, FIMs advance this principle by integrating the 3D flexibility with the large aperture and element density characteristic of metasurfaces. At present, research into FIM-aided communication systems is still in its nascent stages. Initial representative works have begun to establish the potential performance gains across different scenarios. In [10], the joint optimization of transmit beamforming and the FIM surface shape has been investigated to proactively manage inter-user interference and enhance multiuser performance for FIM-aided multiuser systems. Furthermore, the authors in [11] investigated FIM assisted multi-input multi-output (MIMO) systems, where the joint optimization of the FIM 3D surface shape and the transmit covariance matrix was proposed to maximize channel capacity. In [30], the general models for flexible antenna arrays were established and evaluated, which consider various transformations e.g., bending, folding, and rotation, to characterize the resulting channel variations. Beyond enhancing data rates, the unique capabilities of FIMs have been explored for other applications, most notably wireless sensing. the authors in [12] investigated the joint design of FIM surface shapes and transmit waveforms to enhance multi-target sensing performance.
@@ I-B Motivations and Contributions
A common thread throughout the aforementioned application-focused works is the general assumption that the accurate channel is available. However, acquiring the accurate channel in the continuous and high-dimensional deformation space of FIMs is a fundamental and non-trivial challenge [9]. This motivates our work to develop novel and efficient channel estimation techniques specifically tailored for FIM-aided communication systems. Specifically, in contrast to the channel estimation in conventional rigid arrays which involves estimating a finite set of channel parameters for a fixed geometry [22, 23, 24, 25, 26], the channel in an FIM system is a function of the high-dimensional deformation vector that represents the displacement of all radiating elements [10, 11]. Hence, FIM channel estimation requires acquiring a continuous and high-dimensional function that maps the infinite space of all possible physical deformations to the corresponding channel response. The space of possible deformations is continuous and infinite, making it impossible to measure the channel for every possible shape. A naive approach of discretizing this high-dimensional space would lead to an intractable number of pilot transmissions, resulting in prohibitive overhead. Therefore, a core challenge is to develop low-overhead channel estimation schemes that can accurately characterize the channel for the entire continuous deformation space based on a limited number of pilot measurements.
@@ I-C Organizations and Notations
Notations: Lower-case and upper-case boldface letters denote a vector and a matrix, respectively; $\mathbf{A}^{T}$, $\mathbf{A}^{H}$ and $\mathbf{A}^{-1}$ denote the transpose, conjugate transpose and pseudo-inverse of matrix $\mathbf{A}$, respectively; ${\bf{I}}_{a}$ is a $a\times a$ identity matrix; Symbols $|{\cdot}|$, $\left\|{\cdot}\right\|$, and $\left\|{\cdot}\right\|_{F}$ denote the $\ell_{1}$, $\ell_{2}$, and Frobenius norm, respectively; Symbol $\lfloor\cdot\rfloor$ denotes the floor function; Symbol $\odot$ denotes the Hadamard product. Symbol $\propto$ denotes the proportionality relation. Symbol $\langle\cdot,\cdot\rangle$ denotes the inner product between two vectors. Operators $\Re(\mathbf{A})$ and $\Im(\mathbf{A})$ denote the real and imaginary components of the complex-valued $\mathbf{A}$.
II System Model and Problem Formulation
Figure 1: FIM assisted multi-user communications.
As shown in Fig. 1, we consider downlink communication system where a base station (BS) equipped with an FIM serves $K$ single-antenna user equipment (UE). The FIM consists of $N=N_{x}\times N_{z}$ radiating elements arranged as a uniform planar array (UPA) on a flexible substrate. With the aid of controller that can be driven by micro-electromechanical systems (MEMS) or electromagnetic Lorentz force [7, 18, 17], the FIM can programmatically alter its physical shape to optimize the propagation channels for all users.
II-A FIM Geometry and Deformation
@@
[…]
@@
[…]
@@ II-B Channel Model
The channel between the $N$-element FIM at the BS and each of the $K$ UEs is distinct and depends on the deformation shape $\bm{\zeta}$ of FIM. We adopt a clustered multipath channel model, where for each UE $k\in\{1,\dots,K\}$, the channel is formed by $L_{k}$ scattering clusters, and the $l$-th cluster contributes $G_{k,l}$ individual propagation paths. Following the mmWave channel modeling in 3GPP [31], the channel $\mathbf{h}_{k}(\bm{\zeta})\in\mathbb{C}^{N\times 1}$ between BS and UE $k$ is given by
@@
where $\mathbf{o}_{k,l,g}$ is the unit-norm direction of departure (AoD) vector for the $g$-th path in the $l$-th cluster to UE $k$. ${\mathbf{a}}(\bm{\zeta},\mathbf{o})\in\mathbb{C}^{N\times 1}$ is the steering vector of the FIM, which is a function of the deformation $\bm{\zeta}$ and a given AoD $\mathbf{o}$. The AoDs for paths within a single cluster are distributed around a mean cluster angle with a certain angular spread. $\alpha_{k,l,g}$ is the complex gain of the corresponding path, which is defined as [31]
@@
where $d_{k}$ is the distance between the BS and UE $k$. $\beta(d_{k})=P_{0}\left(\frac{d_{k}}{d_{0}}\right)^{-\nu}$ is the large-scale path loss coefficient. Here, $P_{0}$ is the path loss at a reference distance $d_{0}$, and $\nu$ is the path loss exponent. $g_{k,l,g}\sim\mathcal{CN}(0,1)$ is the small-scale fading coefficient, modeled as a complex Gaussian random variable.
In FIM systems, the steering vector of FIM $\mathbf{a}$} can be decomposed into two parts: a rigid component $\mathbf{a}_{\text{rigid}}$ and a flexible component $\mathbf{f}_{\text{flex}}$, i.e.,
@@
$\begin{split}[\mathbf{a}_{\text{rigid}}(\mathbf{o})]_{n}&=e^{j\kappa\langle\mathbf{p}_{n}-\mathbf{p}_{c},\mathbf{o}\rangle}\\ &=e^{j\kappa\langle x_{n}^{t}\mathbf{i}_{t}+y_{n}^{t}\mathbf{j}_{t},\mathbf{o}\rangle},\end{split}$ (7)
@@
$\mathbf{f}_{\text{flex}}(\bm{\zeta},\mathbf{o})]_{n}=e^{j\kappa\langle\zeta_{n}\mathbf{k}_{t},\mathbf{o}\rangle}.$ (8)
In FIM systems, the physical deformation of FIM is typically on the order of a wavelength, e.g., millimeters at mmWave frequencies. Such localized changes are far too small to alter the positions of large-scale scatterers in the environment, e.g., buildings and vehicles. The role of the deformation of FIM is not to change the far-field channel structure but to locally manipulate the phase of the signals at the array by minutely altering the path length to each radiating antenna. This effect is precisely captured in the considered model by the shape-dependent steering vector ${\mathbf{a}}(\bm{\zeta},\mathbf{o})$. Hence, the macroscopic scattering environment, which dictates the path gains and angles, is considered independent of the physical deformation of FIM.
@@
where $\mathbf{u}_{k,l,g}$ denotes the unit-norm direction vector of path (l,g)({l,g}) and can be expressed as
@@
For a UPA-based FIM with a moderate-scale array aperture, the far-field radiation model is adopted due to the fact that the FIM-users distance is greater than the Rayleigh distance [32]. Considering the ideal mutual coupling and pattern distortion for moderate deformations and element spacing near half a wavelength, the array response in FIM systems is given by
$\begin{split}&\left[\mathbf{a}(\bm{\zeta},\theta_{k,l,g},\phi_{k,l,g})\right]_{n}=\\ &e^{j\frac{2\pi}{\lambda}\left(x_{n}\sin\theta_{k,l,g}\cos\phi_{k,l,g}+\zeta_{n}\sin\theta_{k,l,g}\sin\phi_{k,l,g}+z_{n}\cos\theta_{k,l,g}\right)}.\end{split}$ (11)
@@ II-C Problem Formulation
$\begin{split}{\mathbf{y}_{q}}&=\sum\limits_{k=1}^{K}{\mathbf{{h}}_{k}(\bm{\zeta}_{q}){{{{s}}}_{k,q}}+{\mathbf{w}_{q}}},\end{split}$ (12)
@@
Figure 2: Frame protocol for FIM channel estimation.
As illustrated in Fig. 2, suppose the uplink pilot transmission is divided $M$ subframes, in which each subframe is composed of $P$ pilot slots. The feasibility of this frame protocol, which involves transmitting pilots from $M$ distinct shapes within a single channel coherence block, is supported by recent hardware prototypes [17, 18, 8]. In particular, the related prototypes of FIM have demonstrated shape-switching times on the order of milliseconds, which are comparable to or faster than the channel coherence time in typical mmWave scenarios. In the $m$-th subframe, the collected signal $\mathbf{Y}_{m}\in{{\mathbb{C}}^{N\times P}}$ at the BS can be expressed as
@@ III-A Interpolation-Based Channel Estimation
where $(i_{x}(n),i_{z}(n))$ are the 2D indices of the $n$-th element, and $(u,v)$ are spatial frequency indices.
By choosing different pairs of $(u,v)$ and using cosine functions, a rich and structured set of orthogonal basis shapes is generated that effectively probe the channel response to different modes of physical deformation. Compared to random deformation shapes in pilot transmission stage, the superiority of an orthogonal 2D Fourier basis is twofold. First, it guarantees a comprehensive and non-redundant probing of the physical deformation modes of FIMs from low to high spatial frequencies, which is essential for accurately learning the channel function over its continuous domain. Second, this orthogonality systematically minimizes the mutual coherence of the sensing matrix used in the subsequent sparse recovery-based channel estimation algorithms.
@@
In a general $\mathcal{D}$-dimensional space, this concept is extended through barycentric coordinates. First, a simplex, e.g., a triangle in 2D or a tetrahedron in 3D, of $\mathcal{D}+1$ pilot points $\{\bm{\zeta}_{1},\dots,\bm{\zeta}_{\mathcal{D}+1}\}$ that encloses the target shape $\bm{\zeta}_{\text{target}}$ is found via Delaunay triangulation [34]. The target shape can then be uniquely expressed as a linear combination of the simplex vertices:
$\bm{\zeta}_{\text{target}}=\sum_{m=1}^{\mathcal{D}+1}\lambda_{m}\bm{\zeta}_{m}$ (19)
$\text{s.t.}\quad\sum_{m=1}^{\mathcal{D}+1}\lambda_{m}=1\text{ and }\lambda_{m}\geq 0,$
@@
where the coefficients $\{\lambda_{m}\}$ are the barycentric coordinates and serve as the weights in the general interpolation framework. If $\bm{\zeta}_{m}$ is a vertex of the enclosing simplex, $w_{m}(\bm{\zeta}_{\text{target}})=\lambda_{m}$, otherwise $w_{m}(\bm{\zeta}_{\text{target}})=0$. The interpolated channel is then simply $\hat{\mathbf{h}}_{\text{interp}}(\bm{\zeta}_{\text{target}})=\sum_{m=1}^{\mathcal{D}+1}\lambda_{m}\hat{\mathbf{h}}(\bm{\zeta}_{m})$. However, this approach faces significant practical challenges in the high-dimensional space of FIM deformations. For an FIM with $N$ antennas, the deformation space is $N$-dimensional. Constructing a single simplex requires at least $N+1$ pilot shapes. This requirement, $M\geq N+1$, often makes rigorous linear or cubic interpolation computationally infeasible for FIMs where the number of pilot measurements $M$ is typically much smaller than the number of antennas $N$. The infeasibility of performing a rigorous Delaunay triangulation in the high-dimensional FIM deformation space motivates a more practical and local approximation of linear interpolation. Instead of constructing a global triangulation, we propose a method based on finding a local linear representation of the target shape using its KNN from the set of pilot shapes [35]. The KNN-based local linear approach elegantly bypasses the curse of dimensionality associated with global triangulation. By operating in a small, local neighborhood, it remains computationally efficient and numerically stable, while preserving the core principle of linear combination that underlies interpolation methods.